OpenStudy (anonymous):

Solve the equation. 2(y + 7) = 4y + 14

OpenStudy (lgbasallote):

2y + 14 = 4y + 14 -2y = 0 y = 0

OpenStudy (anonymous):

2x-4y=7 (1) 14x-2y=7(2) multiply (1) by -7 -14x+28y =-49 (3) add (2) and (3) 26y= -42 y= -21/13 put this value in (1) 2x+84/13=7---> 26x +84=91 ---> 26x=7 x=7/26

OpenStudy (anonymous):

2x-4y=14x+2y x=-0.5y Then substitute into one of the other equations: 2(-0.5y)-4y=7 y=-7/5 and x=7/10